# Robust Survival Analysis with Adversarial Regularization

### v1

@@ V-A Min-Max Formulation
When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [15]. Typically in SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [16, 6]. Instead, SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of the baseline objective in Section III-B3:
@@
As NNs are highly non-convex, solving this optimization problem is difficult [61]. Therefore, we use the adversarial training methods discussed in Section IV to approximate/relax the inner maximization term of the Min-Max formulation.
@@ V-B Adversarial Robustness Regularization
[21] shows that tight relaxations to the inner maximization of a Min-Max problem may over-regularize the NN, leading to poor predictive performance. Therefore, we extend the baseline objective in Eq. 9 with the adversarial robustness objective in Eq. 19 [55, 57] to balance predictive performance and adversarial robustness as:
$\min_{\boldsymbol{\theta}}\left[\kappa\cdot\mathcal{L}+(1-\kappa)\cdot\max_{\boldsymbol{\tilde{X}}}\mathcal{L}\right],$ (20)
@@
We propose to use the open-source library auto_LiRPA [58] to relax the inner maximization term of the adversarial regularization with CROWN-IBP. auto_LiRPA seamlessly integrates with PyTorch for automatic backpropagation, computing a fully differentiable upper bound on the inner maximization term of Eq. 20 in a computationally feasible manner.
@@ V-C SAWAR Training
We show the pseudo-code in Algorithm 1. The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [62] (12), where we use early stopping [63] with respect to the validation set to prevent overfitting (25, 26 and 27). We introduce the covariate perturbation attacks after 100 epochs (4, 5, 6 and 7). A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window (28) after 100 epochs. We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude is achieved.
Algorithm 1 Robust Training with CROWN-IBP
1:procedure train_step(ModelLoss, Data, $\epsilon$, $\kappa$, Train)
2:   $loss_{epoch}\leftarrow 0$
3:   for $x_{B},t_{B},y_{B}$ in Data do
4:      if $\epsilon>1e^{-20}$ then
5:         lossnatural $\leftarrow\text{ModelLoss}(x_{B},t_{B},e_{B})$
6:         $\tilde{x}_{B}\leftarrow\text{PerturbationLpNorm}(x_{B},\epsilon)$
[…]
8:         loss $\leftarrow\kappa\times loss_{natural}+(1-\kappa)\times loss_{robust}$
9:      else
10:         loss $\leftarrow\text{ModelLoss}(x_{B},t_{B},e_{B})$
11:      if Train then
12:         ADAM Optimizer Step with loss.backward()
13:      lossepoch $\leftarrow$ lossepoch + loss
14:   return lossepoch
15:
16:procedure train_robust(Model, Datatr, Dataval, $\kappa$)
17:   Init random seeds, ADAM optimizer, lr-scheduler, $\epsilon$-scheduler
18:   ModelLoss = Losswrapper(Model)
[…]
20:   for $t$ in Epochs do
21:      if $\epsilon$-scheduler reached $\epsilon_{max}$ then
22:         lr-scheduler Step
[…]
26:         Save Best Model
27:         lossbest $\leftarrow$ lossval
28:      $\epsilon$-scheduler Step
29:   return Model
We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating data uncertainty in SAWAR objective [34] via the adversarial regularization term in Eq. 20 and illustrate SAWAR hypothesis via extensive experiments in the next section.
@@ VI Experiments
We conduct experiments over 10 benchmark SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete. As evaluation metrics, we consider the Concordance Index (CI) [64], the Integrated Brier Score (IBS) [65] and the Negative Log Likelihood (NegLL) at varying $\epsilon$-perturbation magnitudes from 0 to 1. We show that SAWAR improves predictive accuracy, calibration, and population curve estimation.
@@ VI-A Configuration
The experiments are conducted on a 12th Gen Intel(R) Core (TM) 16 GB RAM i9-12900 processor running a Windows operating system. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively. The NN $G_{\boldsymbol{\theta}}$ has two hidden layers of 50 neurons each and Leaky ReLU activation functions. lifelines [66], and sksurv [53] software packages are employed to ensure standardization and reproducibility in computing the SA metrics.
@@ VI-B Datasets
We use the open-source API SurvSet [67], to download the benchmark SA datasets shown in Table I.
Dataset
$N$
$n_{fac}$
$n_{ohe}$
$n_{num}$
TRACE
1878
4
4
2
stagec
146
4
15
3
flchain
7874
4
26
6
Aids2
2839
3
11
1
Framingham
4699
2
12
5
dataDIVAT1
5943
3
14
2
prostate
502
6
16
9
zinc
431
11
18
2
retinopathy
394
5
9
2
LeukSurv
1043
2
24
5
@@
We preprocess each dataset by one-hot encoding categorical covariates, followed by standard normalizing all covariates.
Figure 1: Distribution of survival curves for baseline, noise, FGSM, PGD, and CROWN-IBP. As the strength of the adversarial training method increases, from left to right column, the 90% credible interval narrows. However, the credible interval of SAWAR does not drastically differ from the baseline.
@@ VI-C Adversary Setup
A detailed insight into the robust evaluation procedure is shown in Algorithm 2. We use two adversarial attack methods when benchmarking the robustness of the adversarial regularization method (14, 15 and 19): FGSM attack (Section IV-B) and a novel worst-case attack. We define the worst-case attack as the maximum failure rate with respect to the covariate uncertainty:
$\tilde{\lambda}_{\theta}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (21)
and define the $\epsilon$-worst-case survival curve as:
$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\theta}t},$ (22)
which is the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates. For the exponential distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically. We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq. 21. We evaluate the SA metrics for $\epsilon\in[0,1]$ for the two attacks methods to study how sensitive each adversarial training method is to covariate perturbation (1).
Algorithm 2 Adversarial Robustness Evaluation
1:procedure eval_robust(Model, Data, Metric, Method, $\epsilon$)
2:   $X,t,e$ = Data
3:   MetricValues=[]
4:   for $\epsilon$ in $0:0.1:1$ do
5:      if $\epsilon=0$ then
6:         $\lambda_{\theta}\leftarrow$ Model(X)
7:         MetricValues = [MetricValues, $\text{Metric}(\lambda_{\theta},t,e)$]
8:      else
9:         $\tilde{\lambda}_{\theta}\leftarrow$ ATTACK(Model,Data,Method,$\epsilon$)
10:         MetricValues = [MetricValues, $\text{Metric}(\tilde{\lambda}_{\theta},t,e)$]
11:   return MetricValues
12:
13:procedure attack(Model, Data, Method, $\epsilon$)
14:   $X,t,e$ = Data
15:   if Method = FGSM then
16:      ModelLoss = Losswrapper(Model)
17:      $\tilde{X}\leftarrow$ FGSM(ModelLoss, $X,t,e,\epsilon$)
18:      $\tilde{\lambda}_{\theta}\leftarrow$ Model($\tilde{X}$)
19:   if Method = worst-case then
20:      $\tilde{X}\leftarrow\text{PerturbationLpNorm}(X,\epsilon)$
21:      $\tilde{\lambda}_{\theta}\leftarrow\text{CROWN-IBP}(\text{Model})(\tilde{X},t,e)$
22:   return $\tilde{\lambda}_{\theta}$
VI-D Results
@@
baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR
@@
baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR
@@
Furthermore, we analyze how these adversarial attacks impact the characteristics of the population survival curves.
@@ VI-D1 Failure Rate Distribution
Let us assume that the input covariates are random variables. Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates. Moreover, we may may treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process:
$\boldsymbol{x}\sim p(\boldsymbol{x});\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$ (23)
@@
Furthermore, we compute statistical quantities such as expectation, lower 5th quantile, and upper 95th quantile:
$LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})),$ (25)
$UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})).$ (26)
Fig. 1 visualizes the credible interval and mean of the survival function stochastic process in Eq. 24. The 90% credible interval of the the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 1). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 1). The variance of the failure rate decreasing leads to a more robust model against adversarial attacks. For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$. This is the case for point-estimates of the parameter $\lambda$ derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 20, we trade-off individualized survival curves with vulnerability to adversarial attacks and a population survival curve which is completely resilient to adversarial attacks.
Figure 2: Adversarial robustness of survival curves for baseline, noise, FGSM, PGD, and CROWN-IBP training against $\epsilon$-worst-case attack. As the strength of the adversarial training method increases, from left to right column, the perturbed population survival curve for a given $\epsilon$ is closer to the KMC.
However, despite exhibiting significantly higher adversarial robustness, the credible interval remains relatively unchanged compared to the baseline method.
@@ VI-D2 Population Survival Curve
The $\epsilon$-perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger the adversarial regularization method, the ”closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig. 2). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples $\{t_{i},e_{i}\}_{i=1}^{N}$, to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 2 demonstrates the strong calibration of SAWAR model [29].
We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case attacks (Tables III and II). We empirically showed that CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes.
The exact corresponding metric values for each $\epsilon$ of the worst-case attack are provided in Appendix A.
@@ VII Conclusions
This paper is the first to introduce CROWN-IBP regularization for adversarial robustness training in SA with fully-parametric NN models. We benchmarked SAWAR method on 10 time-to-event datasets with standard adversarial regularization methods such as Gaussian Noise, FGSM, and PGD against the FGSM and worst-case attack. SAWAR’s empirical results indicate that CROWN-IBP regularization consistently enhances performance in SA with respect to CI and IBS, and NegLL for all $\epsilon$-perturbation magnitudes from 0 to 1. Moreover, compared to conventional adversarial regularization methods, SAWAR empirical evidence demonstrates that CROWN-IBP regularization is more resilient to input covariate perturbation attacks, where we see relative improvements to the baseline in CI and IBS of 1%-150% across all $\epsilon$-perturbation magnitudes (Appendix A). Thus adversarial regularization improves predictive performance, adversarial robustness, and calibration.
Future work will include the use the Weibull Proportional Hazard model to enable time varying hazard rates. This will require extending auto_LiRPA to handle the power operator for operations such as $t^{k}$, where $k$ is a learnable parameter.
@@ Appendix A Appendix
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from the baseline to the adversarial training methods becomes larger and then smaller. The relative percent changes from the baseline to SAWAR for $\epsilon$ equal to 1.0, 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05, 0 is 29.46, 48.18, 56.27, 72.23, 102.44, 133.35, 152.17, 140.33, 91.78, 21.17, 9.63, 1.07 respectively. We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
@@
Aids2 baseline 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
@@
TABLE IV: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from the baseline to the adversarial training methods becomes larger and then smaller. The relative percent changes from the baseline to SAWAR for $\epsilon$ equal to 1.0, 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05, 0 is -44.13, -45.46, -46.68, -47.57, -47.86, -47.10, -45.03, -40.89, -33.06, -20.29, -11.34, -1.68 respectively (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
@@
Aids2 baseline 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
@@
TABLE V: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
@@
TABLE VI: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.

### v3

@@ V-A Min-Max Formulation
When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [10]. Typically in SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [11]. Instead, the proposed SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of Eq. 10:
@@
We use the open-source, PyTorch-based library auto_LiRPA [44] to compute the CROWN-IBP upper bound to the inner maximization term in Eq. 20, which is fully differentiable for backpropogation.
@@ V-B Adversarial Robustness Regularization
[41] shows that tight relaxations to the inner maximization of a Min-Max problem can over-regularize the NN, leading to poor predictive performance. Therefore, we balance the objective in Eq. 10 with the adversarial robustness objective in Eq. 20 [40, 43] to combine predictive performance and adversarial robustness as:
$\min_{\boldsymbol{\theta}}\left[\kappa\cdot\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+(1-\kappa)\cdot\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e})\right],$ (21)
@@ V-C SAWAR Training
The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [47], where we use early stopping [48] with respect to the validation set to prevent overfitting We introduce the covariate perturbations after 100 epochs. A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window after 100 epochs. We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude is achieved.
We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating data uncertainty in SAWAR objective [24] via the adversarial regularization term in Eq. 21 and show SAWAR’s improved performance via extensive experiments in the next section.
@@ VI Experiments
We conduct experiments over 10 benchmark medical SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete. As evaluation metrics, we consider the Concordance Index (CI) [49], the Integrated Brier Score (IBS) [50] and the Negative Log Likelihood (NegLL) at varying $\epsilon$-perturbation magnitudes from 0 to 1. To demonstrate that SAWAR produces SA models that perform well and are robust to dataset uncertainty, this section quantifies the improvement in predictive accuracy, calibration, and population curve estimation.
@@ VI-A Configuration
The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively. The NN $G_{\boldsymbol{\theta}}$ has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions. We use lifelines [51] and sksurv [39] for computing SA metrics.
@@ VI-B Datasets
We use the API SurvSet [52] to download the benchmark medical SA datasets shown in Table I and described below:
Dataset $N$ $n_{fac}$ $n_{ohe}$ $n_{num}$
TRACE 1878 4 4 2
stagec 146 4 15 3
flchain 7874 4 26 6
Aids2 2839 3 11 1
Framingham 4699 2 12 5
dataDIVAT1 5943 3 14 2
prostate 502 6 16 9
zinc 431 11 18 2
retinopathy 394 5 9 2
LeukSurv 1043 2 24 5
@@
We preprocess each dataset by one-hot encoding categorical covariates and standard normalization of all covariates.
Figure 1: Distribution of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP. As the strength of the adversarial training method increases, from left to right column, the 90% credible interval narrows. However, the credible interval of SAWAR does not drastically differ from the DRAFT.
Figure 2: Adversarial robustness of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP training against $\epsilon$-worst-case perturbation. As the strength of the adversarial training method increases, from left to right column, the perturbed population survival curve for a given $\epsilon$ is closer to the KMC.
@@ VI-C Adversary Setup
We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B) and a novel worst-case perturbation. FGSM is applied by using the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective Eq. 20, such that:
@@
DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
@@
DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
@@
$\tilde{\lambda}_{\boldsymbol{\theta}}=e^{G_{\boldsymbol{\theta}}\left(\text{FGSM}(\mathcal{L},\boldsymbol{x})\right)}.$ (22)
We define the worst-case perturbation as the maximum failure rate with respect to the covariate uncertainty:
$\tilde{\lambda}_{\boldsymbol{\theta}}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (23)
and the $\epsilon$-worst-case survival curve as the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates:
$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\boldsymbol{\theta}}t}.$ (24)
For the exponential CPH distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically. We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq. 23. We evaluate the SA metrics for $\epsilon\in[0,1]$, where $\epsilon=0$ is no perturbation to input covariates, for the two perturbations methods to study how sensitive each adversarial training method is to covariate perturbation.
VI-D Results
Furthermore, we analyze how adversarial perturbations impact the characteristics of the population survival curves.
@@ VI-D1 Failure Rate Distribution
Let us assume that the input covariates are random variables $\boldsymbol{x}\sim p(\boldsymbol{x})$. Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates $\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$. Moreover, we treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process :
@@
Furthermore, we compute statistical quantities such as expectation, lower 5th quantile $LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$, and upper 95th quantile $UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$.
Fig. 1 visualizes the credible interval and mean of the survival function stochastic process in Eq. 25. The 90% credible interval of the the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 1). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 1). The variance of the failure rate decreasing leads to a more robust model against adversarial perturbations. For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$. This is the case for point-estimates of the parameter $\lambda$ derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 21, we trade-off individualized survival curves with vulnerability to adversarial perturbations and a population survival curve which is completely resilient to adversarial perturbations.
However, despite exhibiting significantly higher adversarial robustness, the credible interval usually remains relatively unchanged compared to the DRAFT method.
@@ VI-D2 Population Survival Curve
The $\epsilon$-perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger the adversarial regularization method, the “closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig. 2). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples $\{t_{i},e_{i}\}_{i=1}^{N}$, to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 2 demonstrates the strong calibration of SAWAR model [19].
We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case perturbations (Tables III and II). We empirically showed that CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes. More detailed results, along with the exact metric values for each $\epsilon$ of the worst-case perturbation for each dataset are available in the code repository linked in the abstract.
@@ VII Conclusions
This paper introduces CROWN-IBP regularization for enhancing adversarial robustness in fully-parametric SA using NN models. We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods such as Gaussian noise, FGSM, and PGD, as well as the novel worst-case perturbation.
Our empirical results show that CROWN-IBP regularization consistently improves performance in SA across various metrics, including CI, IBS, and NegLL, for all $\epsilon$-perturbation magnitudes from 0 to 1. Compared to conventional adversarial regularization techniques, SAWAR demonstrates superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1% to 150% across all $\epsilon$-perturbation magnitudes.
In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration. This approach effectively mitigates data uncertainty in SA and improves generalization across unseen datasets.
Future work will include the use the Weibull CPH model to enable time varying hazard rates. This will require extending auto_LiRPA to handle the power operator for operations such as $t^{k}$, where $k$ is a learnable parameter.
@@ Appendix A Appendix
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in CI from the DRAFT training objective to SAWAR training objective is shown in Table IV (where higher percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
@@
%$\%\Delta$ 29.46 48.18 56.27 72.23 102.44 133.35 152.17 140.33 91.78 21.17 9.63 1.07
TABLE IV: The relative percent change in Concordance Index metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets. A higher relative percent change is better.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 DRAFT 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
@@
TABLE V: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller. The relative percent changes in Integrated Brier Score metric from the DRAFT training objective to SAWAR training objective is shown in Table VI (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
@@
%$\%\Delta$44.13 -45.46 -46.68 -47.57 -47.86 -47.10 -45.03 -40.89 -33.06 -20.29 -11.34 -1.68
TABLE VI: The relative percent change in Integrated Brier Score metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets. A lower relative percent change is better.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 DRAFT 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
@@
TABLE VII: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
@@
TABLE VIII: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.